OpenStudy (anonymous):

tanx=-1.2203 directions: solve the equation on the interval [0,2pie)

myininaya (myininaya):

y=tan(x) has inverse function if we restrict the domain to -pi/2 to pi/2 tan(x)=-1.2203 x=arctan(-1.2203) means we will get an x in between -pi/2 and 0 Use the fact that tan(x) is an an odd function Then you will probably also need to use the fact that tan(x) has period pi

OpenStudy (anonymous):

i just need to know the answer to clarify

myininaya (myininaya):

for example tan(x)=-1 tan(x-pi)=-1 x-pi=arctan(-1) x-pi=-pi/4 x=-pi/4+pi=3pi/4 <--use the fact that tan has period pi You know what we will use that fact again tan(x-2pi)=-1 x-2pi=arctan(-1) x-2pi=-pi/4 x=-pi/4+2pi=7pi/4

myininaya (myininaya):

tan(x-pi*n)=-1.2203 *

myininaya (myininaya):

so same thing here: where you will have one solution for n=1 and the other solution for n=2

OpenStudy (anonymous):

so what are the final solutions @myininaya

myininaya (myininaya):

you will get them from solving the two equations I mentioned

myininaya (myininaya):

solve tan(x-pi*n)=-1.2203 for n=1 solve tan(x-pi*n)=-1.2203 for n=2

OpenStudy (anonymous):

i did but can you give me the solutions so i can see the answers as a solve

myininaya (myininaya):

ok what did you if you solve them?

OpenStudy (anonymous):

what?

myininaya (myininaya):

what did you get when you solved them

OpenStudy (anonymous):

i dont know i dont get it

myininaya (myininaya):

take arctan of both sides

myininaya (myininaya):

just as shown above with that one example

OpenStudy (anonymous):

please give me the solutions i have so much homework left and its 2 in the morning

myininaya (myininaya):

I gave you the set up tan(x-pi*n)=-1.2203 take arctan of both sides... x-pi*n=arctan(-1.2203) now solve for x

myininaya (myininaya):

and evaluate the equation for n=1 and then n=2

OpenStudy (anonymous):

i dont know how

myininaya (myininaya):

that can't be the reason you don't want to do this all I'm asking you to do is add pi*n on both sides and substitute n for 1 then make another equation and substitute n for 2

OpenStudy (anonymous):

i dont know....

OpenStudy (anonymous):

@myininaya please give me the answer